Β Solve the following linear programming problem using the simplex method.

π‘€π‘–π‘›π‘–π‘šπ‘–π‘§π‘’ 𝑃 = 2100𝑦1 + 2400𝑦2 + 10𝑦3 βˆ’ 70𝑦4

𝑠𝑒𝑏𝑗𝑒𝑐𝑑 π‘‘π‘œ

25𝑦1 + 15𝑦2 + 𝑦3 β‰₯ 250

20𝑦1 + 30𝑦2 βˆ’ 𝑦3 βˆ’ 𝑦4 β‰₯ 300

𝑦1 β‰₯ 0, 𝑦2 β‰₯ 0, 𝑦3 β‰₯ 0 , 𝑦4 β‰₯ 0Β 



1
Expert's answer
2021-10-19T10:33:48-0400

Solution:

Given problem is

MinΒ z=2100y1+2400y2+10y3-70y4

subject to

25y1+15y2+y3β‰₯250

20y1+30y2-y3-y4β‰₯300

y1β‰₯0, y2β‰₯0, y3β‰₯0, y4β‰₯0

andΒ y1,y2,y3,y4β‰₯0;


After introducing surplus,artificial variables

MinΒ z=2100y1+2400y2+10y3-70y4+0S1+0S2+0S3+0S4+0S5+0S6+MA1+MA2+MA3+MA4+MA5+MA6

subject to

25y1+15y2+y3-S1+A1=250

20y1+30y2-y3-y4-S2+A2=300

y1-S3+A3=0

y2-S4+A4=0

y3-S5+A5=0

y4-S6+A6=0

andΒ y1,y2,y3,y4,S1,S2,S3,S4,S5,S6,A1,A2,A3,A4,A5,A6β‰₯0



Since allΒ Zj-Cj≀0

Hence, optimal solution is arrived with value of variables as :

y1=0,y2=16.6667,y3=0,y4=200

MinΒ z=26000